# Single-loop Projection-free and Projected Gradient-based Algorithms for Nonconvex-concave Saddle Point Problems with Bilevel Structure

### v1

@@ 2.1 Assumptions and Definitions
Moreover, we impose the following common assumption on the lower-level objective function $g$.
@@
Considering Assumption 2.2-(ii), we can conclude that $\left\|\nabla^{2}_{\theta x}g(x,\theta)\right\|$ is bounded with constant $C^{g}_{\theta x}$ for any $(x,\theta)\in\mathcal{X}\times\mathbb{R}^{m}$.
@@
In this paper, to handle the primal constraint set $\mathcal{X}$ we use two different oracles, namely a linear minimization oracle (LMO) to find $s=\operatorname*{\arg\!\min}_{x\in\mathcal{X}}\left\langle x,\bar{x}\right\rangle$ for any given $\bar{x}\in\mathbb{R}^{n}$ and a projection oracle (PO) to find $s=\operatorname*{\arg\!\min}_{x\in\mathcal{X}}\left\|x-\bar{x}\right\|$ for any given $\bar{x}\in\mathbb{R}^{n}$. Consequently, to measure the quality of solutions generated by the proposed algorithm we define two different gap functions. The first gap function will utilize LMO for the primal part and the generalized gradient norm for the dual part while the second one will use the generalized gradient norm for both the primal and dual parts.
@@
The stationary gap function $\mathcal{G}_{\mathcal{X}}:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$ for the primal part of Problem (1) is defined as $\mathcal{G}_{\mathcal{X}}(x,y)\triangleq\frac{1}{\tau}\left\|s-x\right\|$. The dual gap function in Algorithm 2 is defined similarly to the dual function of Algorithm 1. Additionally, we define $\mathcal{G}_{\mathcal{Z}}(x,y)\triangleq\mathcal{G}_{\mathcal{X}}(x,y)+\mathcal{G}_{\mathcal{Y}}(x,y)$ where $\mathcal{G}_{\mathcal{Y}}$ is introduced in Definition 2.1.
@@
For a given gap function $\mathcal{G}_{\mathcal{Z}}:\mathcal{Z}\rightarrow\mathbb{R}$, where $\mathcal{Z}\triangleq\mathcal{X}\times\mathcal{Y}$, we define a point $(x,y)\in\mathcal{Z}$ as an $\epsilon$-stationary solution for Problem (1) if $\mathcal{G}_{\mathcal{Z}}(x,y)\leq\epsilon$.
@@
$0\in\partial\mathds{1}_{\mathcal{X}}(x)+\nabla_{x}\mathcal{L}(x,y)+u,\quad\text{s.t.}\quad u\in\mathcal{B}_{\mathbb{R}^{n}}(0,\epsilon),$
$0\in\partial\mathds{1}_{\mathcal{Y}}(y)-\nabla_{y}\mathcal{L}(x,y)+v,\quad\text{s.t.}\quad v\in\mathcal{B}_{\mathbb{R}^{d}}(0,\epsilon).$
where we define $\mathcal{B}_{\mathcal{U}}(u,r)\triangleq\big{\{}u\in\mathcal{U}\big{|}\|u\|\leq r\big{\}}$ for a given vector space $\mathcal{U}$.
@@ 2.2 Problem Properties
To solve Problem (1), the initial step involves constructing two implicit gradients for the SP objective function: one corresponding to the minimization variable $x$ and the other for the maximization variable $y$. More specifically,
@@
$\nabla_{y}\mathcal{L}(x,y)=\nabla_{y}\Phi(x,\theta^{*}(x),y)$ (6b)
where ${\bf J}\theta^{*}(x)\in\mathbb{R}^{m\times n}$ is the Jacobian of $\theta^{*}(x)$. One can characterize the Jacobian matrix based on the implicit function theorem for the lower-level problem by writing its first-order optimality condition, namely $\nabla_{\theta}g(x,\theta^{*}(x))=0$ for any $x\in\mathbb{R}^{n}$. By taking the derivative with respect to $x$, one can obtain ${\bf J}\theta^{*}(x)=-\nabla_{\theta x}^{2}g(x,\theta^{*}(x))[\nabla_{\theta\theta}^{2}g(x,\theta^{*}(x))]^{-1}$. Substituting this relation in (6a), we have
$\nabla_{x}\mathcal{L}(x,y)=\nabla_{x}\Phi(x,\theta^{*}(x),y)-\nabla_{\theta x}^{2}g(x,\theta^{*}(x))[\nabla_{\theta\theta}^{2}g(x,\theta^{*}(x))]^{-1}\nabla_{\theta}\Phi(x,\theta^{*}(x),y)$ (7)
@@
Suppose Assumptions 2.1, 2.2 hold. Then for any $x,\overline{x}\in\mathcal{X}$ and $y,\overline{y}\in\mathcal{Y}$, $\left\|v(x,y)-v(\overline{x},\overline{y})\right\|\leq\mathbf{C}_{v1}\left\|x-\overline{x}\right\|+\mathbf{C}_{v2}\left\|y-\overline{y}\right\|$ where $\mathbf{C}_{v1}=\frac{L_{\theta x}^{\Phi}+L_{\theta\theta}^{\Phi}\mathbf{L}_{\theta}}{\mu_{g}}+\frac{C_{\theta}^{\Phi}L_{\theta\theta}^{g}}{\mu^{2}_{g}}(1+\mathbf{L}_{\theta})$ and $\mathbf{C}_{v2}=\frac{1}{\mu_{g}}L_{\theta y}^{\Phi}$.
@@
  1. (I)
    There exists $\mathbf{L}_{\theta}\geq 0$ such that $\left\|\theta^{*}(x)-\theta^{*}(\overline{x})\right\|\leq\mathbf{L}_{\theta}\left\|x-\overline{x}\right\|$ where $\mathbf{L}_{\theta}=\frac{C_{\theta x}^{g}}{\mu_{g}}$.
  2. (II)
    There exists $\mathbf{L}_{\mathcal{L}_{1}},\mathbf{L}_{\mathcal{L}_{2}}\geq 0$ such that $\left\|\nabla_{x}\mathcal{L}(x,y)-\nabla_{x}\mathcal{L}(\overline{x},\overline{y})\right\|\leq\mathbf{L}_{\mathcal{L}_{1}}\left\|x-\overline{x}\right\|+\mathbf{L}_{\mathcal{L}_{2}}\left\|y-\overline{y}\right\|$ where $\mathbf{L}_{\mathcal{L}_{1}}=(L_{xx}^{\Phi}+L_{x\theta}^{\Phi}\mathbf{L}_{\theta}+C_{\theta x}^{g}\mathbf{C}_{v1}+\frac{C_{\theta}^{\Phi}}{\mu_{g}}L_{\theta x}^{g}(1+\mathbf{L}_{\theta})),\mathbf{L}_{\mathcal{L}_{2}}=(L_{xy}^{\Phi}+C_{\theta x}^{g}\mathbf{C}_{v2})$.
  3. (III)
    $\|\nabla_{y}\mathcal{L}(x,y)-\nabla_{y}\mathcal{L}(\overline{x},\overline{y})\|\leq(L_{yx}^{\Phi}+L_{y\theta}^{\Phi}\mathbf{L}_{\theta})\left\|x-\overline{x}\right\|+L_{yy}^{\Phi}\left\|y-\overline{y}\right\|$ where $\mathbf{L}_{\theta}=\frac{C_{\theta x}^{g}}{\mu_{g}}$.
@@ 3 Proposed Methods
In this section, we present our proposed algorithms based on an inexact bilevel regularized primal-dual approach for solving Problem (1). We consider two main approaches to deal with minimization constraint $\mathcal{X}$: the first one involves using a Linear Minimization Oracle (LMO), while the second in the approach we use a Projection Oracle (PO) to update the primal step of the upper-level objective function. To update the maximization variable we consider a gradient ascent-based approach. Moreover, in the lower-level problem, we incorporate one step of gradient-type update to estimate the lower-level optimal solution. As a result, we propose both a one-sided projection-free and a fully projected gradient-based algorithm. Next, we discuss the detailed update of our proposed methods.
@@
$v(x,y)=\operatorname*{\arg\!\min}_{v}\frac{1}{2}v^{\top}\nabla_{\theta\theta}^{2}g(x,\theta^{*}(x))v-\nabla_{\theta}\Phi(x,\theta^{*}(x),y)^{\top}v.$ (9)
@@
$\min_{x\in\mathcal{X}}\max_{y\in\mathcal{Y}}\mathcal{L}_{\mu}(x,y)\triangleq\mathcal{L}(x,y)-\frac{\mu}{2}\|y-y_{0}\|^{2}.$
Moreover, suppose $f_{\mu}:\mathcal{X}\rightarrow\mathbb{R}$ be a function such that $f_{\mu}(x)\triangleq\max_{y\in\mathcal{Y}}\mathcal{L}(x,y)-\frac{\mu}{2}\left\|y-y_{0}\right\|^{2}$, we also could define $y^{*}_{\mu}(x)\triangleq\operatorname*{\arg\!\min}_{y\in\mathcal{Y}}\mathcal{L}(x,y)-\frac{\mu}{2}\left\|y-y_{0}\right\|^{2}$. Here, by controlling $\mu$, we can find $y^{*}_{\mu}(x)$ close to $y^{*}(x)$ (i.e., $y^{*}_{\mu}(x)\approx y^{*}(x)$). Consequently, we can update the dual variable by performing a step of projected gradient ascent, as follows:
$y_{k+1}\leftarrow\mathcal{P}_{\mathcal{Y}}\Big{(}y_{k}+\sigma_{k}\big{(}G_{k}^{y}-\mu(y_{k}-y_{0})\big{)}\Big{)}$
@@
It is noteworthy to mention that our proposed algorithms are designed as single-loop algorithms, making them easy to implement. Moreover, at each iteration, we need at most two Hessian vector products.
@@ 4 Analysis of the Proposed Methods
In this section, we first delve into the analysis of the lower-level approximation in Subsection 4.1. Moving on to Subsection 4.2, we establish a bound on the error between $y_{k}$ and $y_{\mu}^{*}(x_{k},y_{k})$ to scrutinize the trajectory of the optimal solution in the maximization component. Bringing together these subsections, we elucidate the analysis of gradient estimation in Subsection 4.3.
@@ 4.1 Analysis of the Lower-level Approximation
The first step is to analyze the lower-level to see how $\theta_{k}$ is tracking $\theta^{*}(x_{k})$, which is crucial for understanding the convergence behavior of Algorithm 1. This analysis serves to establish bounds on the error between $\theta_{k}$ and $\theta^{*}(x_{k})$, which is essential for assessing the convergence rate of the algorithm.
@@
$\left\|\theta_{k}-\theta^{*}(x_{k})\right\|\leq\beta^{k}\left\|\theta_{0}-\theta^{*}(x_{0})\right\|+\mathbf{L}_{\theta}\sum_{i=0}^{k-1}\beta^{k-i}\left\|x_{i}-x_{i+1}\right\|$ (10)
where $\beta\triangleq(L_{g}-\mu_{g})/(L_{g}+\mu_{g})$.
@@
The inequality in (4.1) together with (12) imply that
@@
$\left\|\theta_{k+1}-\theta^{*}(x_{k+1})\right\|^{2}\leq(\frac{\mu_{g}-L_{g}}{\mu_{g}+L_{g}})^{2}\left\|\theta_{k}-\theta^{*}(x_{k+1})\right\|^{2}$ (14)
@@
Finally, the desired result can be deduced from the above inequality recursively.
In the next step of analyzing the lower-level approximation, we need to understand how well $w_{k+1}$ tracks $v(x_{k},y_{k})$. As explained in Section 3, $v(x_{k},y_{k})$ can be approximated by an iterate $w_{k+1}$ obtained by taking one step of gradient descent. The analysis of the bound on the error between $w_{k+1}$ and $v(x_{k},y_{k})$ is crucial for understanding the behavior and convergence properties of the proposed algorithms.
@@
$\quad+\tilde{C}\left\|\theta_{0}-\theta^{*}(x_{0})\right\|k\rho^{k}+\gamma\tilde{C}\mathbf{L}_{\theta}\sum_{i=0}^{k}\sum_{j=0}^{i-1}\rho^{k-i}\beta^{i-j}\left\|s_{j}-x_{j}\right\|$ (16)
where $\tilde{C}=\eta(L_{\theta\theta}^{g}\frac{C_{\theta}^{\Phi}}{\mu_{g}}+L_{\theta\theta}^{\Phi})$ and $\rho\triangleq(1-\eta\mu_{g})$.
@@
$\quad-\nabla_{\theta\theta}^{2}g(x_{k},\theta^{*}(x_{k}))\Big{)}v(x_{k},y_{k})+\eta\Big{(}\nabla_{\theta}\Phi(x_{k},\theta^{*}(x_{k}),y_{k})-\nabla_{\theta}\Phi(x_{k},\theta_{k},y_{k})\Big{)}\|,$ (17)
@@
$\quad+\eta L_{\theta\theta}^{\Phi}\|\theta_{k}-\theta^{*}(x_{k})\|.$ (18)
@@
$\quad+\eta(L_{\theta\theta}^{g}\frac{C_{\theta}^{\Phi}}{\mu_{g}}+L_{\theta\theta}^{\Phi})\|\theta_{k}-\theta^{*}(x_{k})\|.$ (19)
@@
$\quad+\eta(L_{\theta\theta}^{g}\frac{C_{\theta}^{\Phi}}{\mu_{g}}+L_{\theta\theta}^{\Phi})(\beta^{i}\left\|\theta_{0}-\theta^{*}(x_{0})\right\|+\mathbf{L}_{\theta}\sum_{j=0}^{i-1}\beta^{i-j}\left\|x_{j}-x_{j+1}\right\|).$ (20)
@@
$\quad+\tilde{C}\left\|\theta_{0}-\theta^{*}(x_{0})\right\|k\rho^{k}+\tilde{C}\mathbf{L}_{\theta}\sum_{i=0}^{k}\sum_{j=0}^{i-1}\rho^{k-i}\beta^{i-j}\left\|x_{j}-x_{j+1}\right\|$

### v3

@@ 2.1 Assumptions and Definitions
Moreover, we impose the following common assumptions on the lower-level objective function $g$.
@@
Considering Assumption 2.2-(2), we can conclude that $\left\|\nabla^{2}_{\theta x}g(x,\theta)\right\|$ is bounded with constant $C^{g}_{\theta x}$ for any $(x,\theta)\in\mathcal{X}\times\mathbb{R}^{m}$.
@@
In this paper, to handle the primal constraint set $\mathcal{X}$, we use two different oracles, namely a linear minimization oracle (LMO) to find $s=\operatorname*{\arg\!\min}_{x\in\mathcal{X}}\left\langle x,\bar{x}\right\rangle$ for any given $\bar{x}\in\mathbb{R}^{n}$ and a projection oracle (PO) to find $s=\operatorname*{\arg\!\min}_{x\in\mathcal{X}}\left\|x-\bar{x}\right\|$ for any given $\bar{x}\in\mathbb{R}^{n}$. Consequently, to measure the quality of solutions generated by the proposed algorithms, we define two different gap functions. The first gap function will utilize LMO for the primal part and the generalized gradient norm for the dual part while the second one will use the generalized gradient norm for both the primal and dual parts.
@@
The stationary gap function $\mathcal{G}_{\mathcal{X}}:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$ for the primal part of Problem (1) is defined as $\mathcal{G}_{\mathcal{X}}(x,y)\triangleq\frac{1}{\tau}\left\|x-\mathcal{P}_{\mathcal{X}}(x-\tau\nabla_{x}\mathcal{L}(x,y))\right\|$. The dual gap function is defined similarly to Definition 2.1. Additionally, we define $\mathcal{G}_{\mathcal{Z}}(x,y)\triangleq\mathcal{G}_{\mathcal{X}}(x,y)+\mathcal{G}_{\mathcal{Y}}(x,y)$ where $\mathcal{G}_{\mathcal{Y}}$ is introduced in Definition 2.1.
@@
For a given gap function $\mathcal{G}_{\mathcal{Z}}:\mathcal{Z}\rightarrow\mathbb{R}$, where $\mathcal{Z}\triangleq\mathcal{X}\times\mathcal{Y}$, we call a point $(x,y)\in\mathcal{Z}$ an $\epsilon$-stationary solution for Problem (1) if $\mathcal{G}_{\mathcal{Z}}(x,y)\leq\epsilon$.
@@
$0\in\partial\mathds{1}_{\mathcal{X}}(x)+\nabla_{x}\mathcal{L}(x,y)+u,\quad\text{s.t.}\quad u\in\mathcal{B}_{\mathbb{R}^{n}}(0,\epsilon)$
$0\in\partial\mathds{1}_{\mathcal{Y}}(y)-\nabla_{y}\mathcal{L}(x,y)+v,\quad\text{s.t.}\quad v\in\mathcal{B}_{\mathbb{R}^{d}}(0,\epsilon),$
where we define $\mathcal{B}_{\mathcal{U}}(\bar{u},r)\triangleq\big{\{}u\in\mathcal{U}\big{|}\|u-\bar{u}\|\leq r\big{\}}$ for a given vector space $\mathcal{U}$ and $\bar{u}\in\mathcal{U}$.
@@ 2.2 Problem Properties
Problem (1) can be viewed as an implicit SP problem, i.e., $\min_{x\in\mathcal{X}}\max_{y\in\mathcal{Y}}\mathcal{L}(x,y)$. Therefore, it is natural to first investigate the two implicit partial gradients of the objective function: one corresponding to the minimization variable $x$ and the other to the maximization variable $y$. More specifically, due to the uniqueness of the solution and smoothness of the lower-level objective function one can show that
@@
$\nabla_{y}\mathcal{L}(x,y)=\nabla_{y}\Phi(x,\theta^{*}(x),y),$ (3b)
where ${\bf J}\theta^{*}(x)\in\mathbb{R}^{m\times n}$ is the Jacobian of $\theta^{*}(x)$. The Jacobian matrix can be characterized based on the implicit function theorem for the lower-level problem by writing its first-order optimality condition, namely $\nabla_{\theta}g(x,\theta^{*}(x))=0$ for any $x\in\mathbb{R}^{n}$. By taking the derivative with respect to $x$, one can obtain ${\bf J}\theta^{*}(x)=-\nabla_{\theta x}^{2}g(x,\theta^{*}(x))[\nabla_{\theta\theta}^{2}g(x,\theta^{*}(x))]^{-1}$. Substituting this relation in (3a), we have
$\nabla_{x}\mathcal{L}(x,y)=\nabla_{x}\Phi(x,\theta^{*}(x),y)-\nabla_{\theta x}^{2}g(x,\theta^{*}(x))[\nabla_{\theta\theta}^{2}g(x,\theta^{*}(x))]^{-1}\nabla_{\theta}\Phi(x,\theta^{*}(x),y).$
@@
Suppose Assumptions 2.1 and 2.2 hold. Then for any $x,\overline{x}\in\mathcal{X}$ and $y,\overline{y}\in\mathcal{Y}$, we have that $\left\|v(x,y)-v(\overline{x},\overline{y})\right\|\leq\mathbf{C}_{v1}\left\|x-\overline{x}\right\|+\mathbf{C}_{v2}\left\|y-\overline{y}\right\|$ where $\mathbf{C}_{v1}\triangleq\frac{L_{\theta x}^{\Phi}+L_{\theta\theta}^{\Phi}\mathbf{L}_{\theta}}{\mu_{g}}+\frac{C_{\theta}^{\Phi}L_{\theta\theta}^{g}}{\mu^{2}_{g}}(1+\mathbf{L}_{\theta})$ and $\mathbf{C}_{v2}\triangleq\frac{1}{\mu_{g}}L_{\theta y}^{\Phi}$.
Proof.
See Appendix 8.1 for the proof.
@@
(I) There exists $\mathbf{L}_{\theta}\geq 0$ such that $\left\|\theta^{*}(x)-\theta^{*}(\overline{x})\right\|\leq\mathbf{L}_{\theta}\left\|x-\overline{x}\right\|$ where $\mathbf{L}_{\theta}\triangleq\frac{C_{\theta x}^{g}}{\mu_{g}}$.
(II) There exists $\mathbf{L}_{\mathcal{L}_{1}},\mathbf{L}_{\mathcal{L}_{2}}\geq 0$ such that $\left\|\nabla_{x}\mathcal{L}(x,y)-\nabla_{x}\mathcal{L}(\overline{x},\overline{y})\right\|\leq\mathbf{L}_{\mathcal{L}_{1}}\left\|x-\overline{x}\right\|+\mathbf{L}_{\mathcal{L}_{2}}\left\|y-\overline{y}\right\|$ where $\mathbf{L}_{\mathcal{L}_{1}}\triangleq(L_{xx}^{\Phi}+L_{x\theta}^{\Phi}\mathbf{L}_{\theta}+C_{\theta x}^{g}\mathbf{C}_{v1}+\frac{C_{\theta}^{\Phi}}{\mu_{g}}L_{\theta x}^{g}(1+\mathbf{L}_{\theta})),\mathbf{L}_{\mathcal{L}_{2}}\triangleq(L_{xy}^{\Phi}+C_{\theta x}^{g}\mathbf{C}_{v2})$.
(III) $\|\nabla_{y}\mathcal{L}(x,y)-\nabla_{y}\mathcal{L}(\overline{x},\overline{y})\|\leq(L_{yx}^{\Phi}+L_{y\theta}^{\Phi}\mathbf{L}_{\theta})\left\|x-\overline{x}\right\|+L_{yy}^{\Phi}\left\|y-\overline{y}\right\|$.
Proof.
See Appendix 8.2 for the proof.
@@ 3 Proposed Methods
In this section, we present our proposed algorithms based on an inexact bilevel regularized primal-dual approach for solving Problem (1). We consider two main approaches to deal with the constraint $\mathcal{X}$ in the minimization: the first one involves using a Linear Minimization Oracle (LMO), while in the second approach, we use a Projection Oracle (PO) to update the primal step of the upper-level objective function. To update the maximization variable, we consider a gradient ascent-based approach. Moreover, in the lower-level problem, we incorporate one step of gradient-type update to estimate the lower-level optimal solution. As a result, we propose both a one-sided projection-free and a fully projected gradient-based algorithms. Next, we discuss the detailed update of our proposed methods.
@@
$v(x,y)=\operatorname*{\arg\!\min}_{v}\frac{1}{2}v^{\top}\nabla_{\theta\theta}^{2}g(x,\theta^{*}(x))v-\nabla_{\theta}\Phi(x,\theta^{*}(x),y)^{\top}v.$
@@
$\min_{x\in\mathcal{X}}\max_{y\in\mathcal{Y}}\mathcal{L}_{\mu}(x,y)\triangleq\mathcal{L}(x,y)-\frac{\mu}{2}\|y-y_{0}\|^{2}.$ (5)
Moreover, suppose $f_{\mu}:\mathcal{X}\rightarrow\mathbb{R}$ be a function such that $f_{\mu}(x)\triangleq\max_{y\in\mathcal{Y}}\mathcal{L}(x,y)-\frac{\mu}{2}\left\|y-y_{0}\right\|^{2}$, we also define $y^{*}_{\mu}(x)\triangleq\operatorname*{\arg\!\min}_{y\in\mathcal{Y}}\mathcal{L}(x,y)-\frac{\mu}{2}\left\|y-y_{0}\right\|^{2}$. Here, by controlling $\mu$, we can find $y^{*}_{\mu}(x)$ close to $y^{*}(x)$ (i.e., $y^{*}_{\mu}(x)\approx y^{*}(x)$). Consequently, we can update the dual variable by performing a step of projected gradient ascent, as follows:
$y_{k+1}\leftarrow\mathcal{P}_{\mathcal{Y}}\Big{(}y_{k}+\sigma_{k}\big{(}G_{k}^{y}-\mu(y_{k}-y_{0})\big{)}\Big{)}.$
@@
It is noteworthy to mention that our proposed algorithms are designed as single-loop algorithms requiring at most two Hessian vector products.
@@ 4 Analysis of the Proposed Methods
In this section, we first delve into the analysis of the lower-level approximation in Subsection 4.1. Moving on to Subsection 4.2, we establish a bound on the error between $y_{k}$ and $y_{\mu}^{*}(x_{k})$ to examine the trajectory of the optimal solution in the maximization component. Bringing together these subsections, we elucidate the analysis of gradient estimation in Subsection 4.3.
@@ 4.1 Analysis of the Lower-level Approximation
The first step is to analyze the lower-level to see how $\theta_{k}$ is tracking $\theta^{*}(x_{k})$, which is crucial for understanding the convergence behavior of proposed algorithms. This analysis serves to establish a bound on the error between $\theta_{k}$ and $\theta^{*}(x_{k})$, which is essential for assessing the convergence rate of the algorithm.
@@
$\left\|\theta_{k}-\theta^{*}(x_{k})\right\|\leq\beta^{k}\left\|\theta_{0}-\theta^{*}(x_{0})\right\|+\mathbf{L}_{\theta}\sum_{i=0}^{k-1}\beta^{k-i}\left\|x_{i}-x_{i+1}\right\|,$
where $\beta\triangleq(L_{g}-\mu_{g})/(L_{g}+\mu_{g})$. Moreover, the sequence $\{\theta_{k}\}_{k\geq 0}$ is a bounded sequence.
@@
The inequality in (7) together with (4.1) imply that
@@
$\left\|\theta_{k+1}-\theta^{*}(x_{k+1})\right\|^{2}\leq(\frac{\mu_{g}-L_{g}}{\mu_{g}+L_{g}})^{2}\left\|\theta_{k}-\theta^{*}(x_{k+1})\right\|^{2}.$
@@
Therefore, the desired result can be deduced from the above inequality recursively.
Moreover, since $\mathcal{X}$ is compact, the sequence $\{x_{k}\}$ remains in a bounded set, implying that $\|x_{i}-x_{i+1}\|$ is uniformly bounded for all $i$. Given that $\theta^{*}(\cdot)$ is a continuous map on the compact set $\mathcal{X}$, it is also bounded. Consequently, since $10<\beta<$ 1, the right-hand side of the inequality remains bounded, ensuring that the sequence $\{\theta_{k}\}_{k\geq 0}$ is a bounded sequence.
In the next step of analyzing the lower-level approximation, we need to understand how well $w_{k+1}$ tracks $v(x_{k},y_{k})$. As explained in Section 3, $v(x_{k},y_{k})$ can be approximated by an iterate $w_{k+1}$ obtained by taking one step of gradient descent.
@@
$\quad+\tilde{C}\left\|\theta_{0}-\theta^{*}(x_{0})\right\|k\rho^{k}+\gamma\tilde{C}\mathbf{L}_{\theta}\sum_{i=0}^{k}\sum_{j=0}^{i-1}\rho^{k-i}\beta^{i-j}\left\|s_{j}-x_{j}\right\|,$
where $\tilde{C}=\eta(L_{\theta\theta}^{g}\frac{C_{\theta}^{\Phi}}{\mu_{g}}+L_{\theta\theta}^{\Phi})$ and $\rho\triangleq(1-\eta\mu_{g})$. Moreover, the sequence $\{w_{k}\}_{k\geq 0}$ is a bounded sequence.
@@
$\quad-\nabla_{\theta\theta}^{2}g(x_{k},\theta^{*}(x_{k}))\Big{)}v(x_{k},y_{k})+\eta\Big{(}\nabla_{\theta}\Phi(x_{k},\theta^{*}(x_{k}),y_{k})-\nabla_{\theta}\Phi(x_{k},\theta_{k},y_{k})\Big{)}\|,$
@@
$\quad+\eta L_{\theta\theta}^{\Phi}\|\theta_{k}-\theta^{*}(x_{k})\|.$
@@
$\quad+\eta(L_{\theta\theta}^{g}\frac{C_{\theta}^{\Phi}}{\mu_{g}}+L_{\theta\theta}^{\Phi})\|\theta_{k}-\theta^{*}(x_{k})\|.$
@@
$\quad+\eta(L_{\theta\theta}^{g}\frac{C_{\theta}^{\Phi}}{\mu_{g}}+L_{\theta\theta}^{\Phi})(\beta^{i}\left\|\theta_{0}-\theta^{*}(x_{0})\right\|+\mathbf{L}_{\theta}\sum_{j=0}^{i-1}\beta^{i-j}\left\|x_{j}-x_{j+1}\right\|).$
@@
$\quad+\tilde{C}\left\|\theta_{0}-\theta^{*}(x_{0})\right\|k\rho^{k}+\tilde{C}\mathbf{L}_{\theta}\sum_{i=0}^{k}\sum_{j=0}^{i-1}\rho^{k-i}\beta^{i-j}\left\|x_{j}-x_{j+1}\right\|,$
@@
Moreover, since $\mathcal{X}$ and $\mathcal{Y}$ are compact, and $v(\cdot,\cdot)$ is a continuous map on these sets, it follows that $v$ is bounded. Additionally, all terms on the right-hand side are bounded implying that the sequence $\{w_{k}\}_{k\geq 0}$ is also bounded.